You have gathered the following information on your investments. What is the expected return on the portfolio?

Stock Number of Shares Price per Share Expected Return
F 190 $ 28 12.84 %
G 255 $ 14 9.45 %
H 195 $ 40 10.23 %
Answer #1
Stock Total value=Number of shares*Price/Share Expected Return
F (190*28)=5320 12.84%
G (255*14)=3570 9.45%
H (195*40)=7800 10.23%
Total=$16690

Portfolio return=Respective return*Respective investment weight

=(5320/16690*12.84)+(3570/16690*9.45)+(7800/16690*10.23)

=10.9%(Approx).